A random sample is drawn from a population of unknown standard deviation. Construct a 99% confidence interval for the population mean based on the information given:

a. N=49 ×=17.1 0=2.1

b. n=169 ×=17.1 0=2.1


1
Expert's answer
2021-03-09T08:08:49-0500

a. The provided sample mean is "\bar{x}=17.1"  and the sample standard deviation is "s=2.1." The size of the sample is "n=49" and the required confidence level is "99\%."

The number of degrees of freedom are "df=n-1=49-1=48," and the significance level is "\alpha=0.01."

Based on the provided information, the critical "t" -value for "\alpha=0.01" and "df=48" degrees of freedom is "t_c=2.682204."

The 99% confidence for the population mean "\mu"  is computed using the following expression


"CI=(\bar{x}-\dfrac{t_c\times s}{\sqrt{n}},\bar{x}+\dfrac{t_c\times s}{\sqrt{n}} )"

"CI=(17.1-\dfrac{2.682204\times 2.1}{\sqrt{49}},17.1-\dfrac{2.682204\times 2.1}{\sqrt{49}} )"

"=(16.2953, 17.9047)"

b. The provided sample mean is "\bar{x}=17.1"  and the sample standard deviation is "s=2.1." The size of the sample is "n=169" and the required confidence level is "99\%."

The number of degrees of freedom are "df=n-1=169-1=168," and the significance level is "\alpha=0.01."

Based on the provided information, the critical "t" -value for "\alpha=0.01" and "df=168" degrees of freedom is "t_c=2.60541."

The 99% confidence for the population mean "\mu"  is computed using the following expression


"CI=(\bar{x}-\dfrac{t_c\times s}{\sqrt{n}},\bar{x}+\dfrac{t_c\times s}{\sqrt{n}} )"

"CI=(17.1-\dfrac{2.60541\times 2.1}{\sqrt{169}},17.1-\dfrac{2.60541\times 2.1}{\sqrt{169}} )"

"=(16.6791, 17.5209)"